OpenStudy (anonymous):

car traveling 95 km/h is 110 m behind a truck traveling 75 km/h.How long will it take the car to reach truck? i want to know which equation i should use and explain how to answer it step by step

OpenStudy (raffle_snaffle):

Lets first draw a picture.

OpenStudy (raffle_snaffle):

Can you draw me a picture of what is being demonstrated in the problem?

OpenStudy (anonymous):

well its question without picture , i cant sorry

OpenStudy (raffle_snaffle):

You have to draw the picture like in most physics problems

OpenStudy (raffle_snaffle):

Which kinematic equations do you think we use and why?

OpenStudy (anonymous):

 use meter because the distance between car and truck by meter

OpenStudy (snowsurf):

Before plugging values into equations you need to convert the units to meters per second. I will leave that you to do. For the truck \[x = x_0 + v_Tt\] For car \[x = x_0 + v_Ct\] Let pick a reference frame to see how this will work. I will sue Cartesian coordinate system. We have the car placed at the origin and truck is 110 meter from the origin. We know that truck is still moving so that means it still increase distance along the x axis. Sorry I forgot to label the coordinates. When know when the car catches up to the truck their distances will be the same. Since the car is at the origin we can say its initial position \[x_0 = 0\] So for the car we have \[x = v_Ct\] And for the truck we have \[x = x_0 + v_Tt\] When the car catches up to the truck distances will be the same. So we have \[x_0 + v_Tt = v_Ct\] Now solve for t \[\frac{ x_0}{ (v_C-v_T) } = t\] Where \[v_c\] is the velocity of the car and \[v_T\] is the velocity of the truck.